OpenStudy (anonymous):

Help? If the perimeter of triangle ABC is 50 cm, what is the perimeter of triangle DEF?

OpenStudy (anonymous):

 ABC~DEF

OpenStudy (amistre64):

all linear aspects are scaled, so whatever the scale ratio is you multiply by

OpenStudy (amistre64):

spose we have sides: a,b,c P = a+b+c and some scale factor n such that P' = an+bn+cn = n(a+b+c)

OpenStudy (anonymous):

scale factor=21/15=7/5 p=(7/5)*50=?

OpenStudy (anonymous):

70

OpenStudy (anonymous):

Could you help me with one more problem? @surjithayer